# A Survey on All-in-One Image Restoration: Taxonomy, Evaluation and Future Trends

### v1

@@ Abstract
Image restoration (IR) to the process of improving visual quality of images while removing degradation, such as noise, blur, weather effects, and so on. Traditional IR methods typically target specific types of degradation, which limits their effectiveness in real-world scenarios with complex distortions. In response to this challenge, the all-in-one image restoration (AiOIR) paradigm has emerged, offering a unified framework that adeptly addresses multiple degradation types. These innovative models enhance both convenience and versatility by adaptively learning degradation-specific features while simultaneously leveraging shared knowledge across diverse corruptions. In this review, we delve into the AiOIR methodologies, emphasizing their architecture innovations and learning paradigm and offering a systematic review of prevalent approaches. We systematically categorize prevalent approaches and critically assess the challenges these models encounter, proposing future research directions to advance this dynamic field. Our paper begins with an introduction to the foundational concepts of AiOIR models, followed by a categorization of cutting-edge designs based on factors such as prior knowledge and generalization capability. Next, we highlight key advancements in AiOIR, aiming to inspire further inquiry and innovation within the community. To facilitate a robust evaluation of existing methods, we collate and summarize commonly used datasets, implementation details, and evaluation metrics. Additionally, we present an objective comparison of open-sourced methods, providing valuable insights for researchers and practitioners alike. This paper stands as the first comprehensive and insightful review of AiOIR. A related repository is available at https://github.com/Harbinzzy/All-in-One-Image-Restoration-Survey.
@@ I Introduction
Image processing is an integral part of low-level vision tasks, and digital image processing has evolved significantly over the past few decades, transitioning from traditional methods to advanced deep learning techniques. Initially, image processing relied heavily on algorithms that performed tasks such as filtering, edge detection, image synthesis and image segmentation. These methods, while effective, were limited by their inability to handle complex and varied image degradation scenarios. With the rise of deep learning, image processing has achieved remarkable results, especially driven by convolutional neural networks (CNNs) [1], Transformer [2] and StableDiffusion [3]. Among the image restoration field, single-task IR have achieved notable breakthroughs, which focuses on correcting a specific type of image degradation(e.g., denoising [4, 5, 6, 7, 8, 9], dehazing [10, 11, 12, 13, 14], desnowing [15, 16, 17], deraining [18, 19, 20], deblurring [21, 22] and low-light image enhancement [23, 24, 25]). Although they have set state-of-the-art performance on restoring diverse degraded images, existing single-task IR methods: 1) lack the flexibility to adapt to new types of image degradations without extensive retraining; 2) typically require separate models for each type of degradation, which can be resource-intensive and impractical in real-world applications where images often suffer from multiple types of degradations simultaneously; 3) employ fixed network structures, which restrict the diversity of corruption-specific knowledge. To alleviate the aforementioned issues, researchers have proposed all-in-one image restoration (AiOIR), devoting significant effort to introduce various key improvements(e.g., prompt-based learning approach, Mixture-of-Experts (MoE) structure and multimodal model) beyond single-task IR.
The AiOIR approach aims to develop models capable of handling multiple degradations tasks simultaneously within a unified framework. AiOIR integrates a variety of image enhancement and restoration techniques into a single model, offering greater performance and generalizability. These models are more efficient, as they eliminate the need for multiple specialized models, and they are more robust, as they can adapt to different types of degradations presented in an image. It is worth noting that diverse and unforeseen degradations may arise in the real-world scenarios jointly, hence the need for AiOIR models further developing. In some works, the corruption type label is available while inferencing, namely non-blind IR. In contrast, if the input images are without priors, which is called blind IR. If we focus on whether the training set and the test set are the same, that is, whether it’s zero-shot, we can divide the tasks into open-set scenario IR and close-set scenario IR. Despite the variety of AiOIR models, their ability to generate high-quality images is being actively explored. Recently, pioneering researchers have been investigating ways to optimize the architecture of these models to balance the trade-offs between computational complexity and restoration quality. We have listed the representative works on all-in-one models based on the timeline in Fig. 2.
@@
With the rapid development of AiOIR, numerous researchers collected a series of datasets tailored for various IR tasks, e.g., BSD100 [26], Manga109 [27] and Urban100 [28] for SR, RainDrop [29], Outdoor-Rain [30], SPA [31] and Rain-100H [32] for draining, GoPro [33] and HIDE [34] for deblurring, LOL [35] for low-light image enhancement, etc. Leveraging these datasets, the majority of recent works focused on improving the representation capability of IR networks for complicated degradation through well-designed methods based on prompt learning, contrast learning, multimodel representations, etc. Although these works achieve superior progress in the objective quality(e.g., PSNR, SSIM [36], NIQE [37] and FID [38]), the restored images still suffer from unsatisfied texture generation, hindering the application of IR methods in real-world scenarios. Collectively, these methodologies represent a substantial progression in the pursuit of sophisticated, accurate, and versatile AiOIR solutions.
@@
AiOIR methods have emerged as a positive step forward, although the details and applications of these approaches still need to be worked out. Keeping up with the latest research is vital for CV researchers. While there are comprehensive surveys for single-task IR methods, such as image super-resolution [39, 40, 41, 42], deraining [43, 44], dehazing [45], denoising [46, 47, 48], deblurring [49, 50], low-light enhancement [51, 52, 53] and reviews covering diffusion-based image restoration [54], there is currently no review that specifically addresses the AiOIR field. This gap limits its development. Consequently, we aim to provide the first comprehensive overview of AiOIR methods in IR tasks, shedding light on both its representative approaches and diverse improvements. Fig. 1 shows the taxonomy of this survey in a hierarchically-structured way. In the following sections, we will cover various aspects of recent advances in AiOIR:
@@
    Preliminaries (Sec. II) : We present the concept definition of AiOIR and related concepts comparison. Meanwhile, We point out different task settings and typical scenarios.
@@
    Methodology (Sec. III) : We provide a detailed analysis of several representative AiOIR networks, aiming to illustrate the prevailing methods and elucidate the different categories of methods. By analyzing the methods used to achieve state-of-the-art, we summarized network architectures, basic methods, learning strategy as well as some key improvements.
@@
    Experiments (Sec. IV) : To facilitate a reasonable and exhaustive comparison, we clarify the commonly-used datasets, and experimental settings in different AiOIR tasks. Furthermore, the comprehensive comparisons between the benchmarks across different tasks are provided.
@@
    Challenges and Future Directions (Sec. V) : There are still some challenges to extending them to the practical application. To further improve the development of AiOIR, we summarize the primary challenges and propose the potential directions and trends for solving them.
Figure 2: Representative works in AiOIR based on the timeline. The bar chart shows number of papers in AiOIR from 2019 to October 2024.
@@ II Preliminaries
In this section, we first define the concept of All-in-One image restoration and compare it with related concepts. Subsequently, we examine task settings in AiOIR from various perspectives. Finally, we provide a comprehensive review of diverse scenarios within AiOIR and the corresponding approaches.
II-A Concept Definition
@@
Single degradation image restoration (SDIR) focuses on recovering clean images from observations that have been corrupted by a specific type of degradation, such as noise, blur, or rain. These methods are typically task-specific, meaning that each model is designed for a particular degradation type. While these models excel in handling known degradation types, they often struggle when faced with unseen degradations or new levels of corruption. Consequently, multiple degradation image restoration (MDIR) addresses the challenge of processing images affected by various degradation types. The primary strategy involves training a single model using datasets encompassing diverse degradation types. Although the model encounters multiple tasks during training, it lacks specialized design elements to effectively manage the coexistence of different degradations. As a result, MDIR models frequently exhibit mediocre performance across tasks due to their inability to provide a true all-in-one solution specifically. To tackle this issue, researchers propose AiOIR, which aims to recover clean images under various degraded conditions explicitly tailored to address multiple degradations based on multi-head and multi-tail structures, prior, or pretrained models within an unified framework. These models present practical advantages such as reduced storage requirements and simplified deployment; however, the principal challenge lies in developing robust architectures capable of effectively addressing diverse degradations with a singular set of parameters while achieving high-quality restorations across varying conditions. In addition, these approaches often necessitate prior knowledge regarding the type or extent of degradation involved limiting their applicability in real-world scenarios where such information is unknown or variable. In subsequent sections, we will elucidate key concepts related to degradation awareness and generalization capabilities pertinent to all-in-one models.
II-B Task Settings
II-B1 Blind/Non-Blind
We can preliminarily categorize AiOIR methods into two main types: blind and non-blind, based on whether the type or level of degradation information is known or not. Non-blind image restoration assumes that the degradation process is either known or can be explicitly modeled. For instance, when the blurring kernel or noise distribution is available, the restoration task primarily focuses on recovering the original image based on this known information. As a result, non-blind restoration is typically considered less challenging than blind restoration, as it involves fewer unknowns. Some works [55, 56, 57, 58, 59, 60, 61, 62] require prior knowledge of the degradation in the corrupted image to feed it to the appropriate restoration head or guide the design of the image restoration network. It is worthy to note that a more challenging scenario is the opposite case [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78], where the degradation affecting the input image is unknown—this defines blind AiOIR, which has shown promising potential for real-world photographs.
In blind AiOIR, the degradation type (e.g., blurring, noise, or compression) is unknown and should be estimated concurrently with the recovery of the original image. The difficulty in blind restoration arises from this dual estimation task, which requires advanced techniques to effectively model both the degradation process and the image restoration simultaneously. As a result, blind restoration often demands more sophisticated algorithms or more complex models to estimate the reconstructed image accurately. To achieve blind AiOIR, AirNet [63] learns degradation representations from corrupted images using a contrastive learning strategy. These learned representations are then employed to restore the clean image. Subsequently, IDR [66] models various degradations based on underlying physical principles and accomplishes AiOIR in two stages. More recently, several prompt-learning-based approaches have emerged [79, 80, 61, 65, 81, 82, 83, 79, 84]. For instance, PromptIR [65] introduces a series of learnable prompts to encode discriminative information about different degradation types, involving a substantial number of parameters.
II-B2 Open-set/Closed-set
In the field of AiOIR, a novel distinction exists between close-set and open-set image restoration, which addresses the consistency and generalization between training and testing scenarios. Close-set image restoration assumes a predefined set of degradations encountered during the training phase, restricting the model’s application to scenarios where the degradations in test images are identical to those in the training data. This approach, while effective for known degradations (i.e., close-set scenario), lacks the flexibility to handle unexpected or unseen types of degradation, which are more common in real-world applications.
In contrast, open-set image restoration (e.g., Gou et al. [85]) addresses the challenge of restoring images affected by unforeseen degradations that are absent from the training data. Meanwhile, unlike zero-shot IR, which focuses on the recovery of a single degraded image without relying on pre-existing degradation-specific training. Although free from the constraints of training on predefined degradations, zero-shot IR methods often require some form of prior knowledge or assumptions about the test degradations in advance to guide the restoration process. The core difficulty in open-set scenarios lies in the distribution shifts between training and test data. This requires models to generalize beyond their training experience and adapt to previously unseen degradation types. Test-time adaptation (TTA) (e.g.[86, 87, 88]) has emerged as a key technique in addressing this issue, allowing models to dynamically adjust their parameters during the testing phase to better align with the characteristics of the degraded input images. TTA effectively addresses distribution shifts between test and training data by adjusting the pre-trained model using test samples. By enabling the model to restore images with unknown degradations, open-set image restoration aims to provide a more robust and versatile solution, making it a crucial step towards real-world applicability in diverse image restoration tasks.
II-B3 Zero-shot
Zero-shot image restoration involves restoring images with distortions not encountered during training, requiring the model to classify unknown classes into specific categories, unlike Open-set methods [89]. This approach demands a robust and adaptable model, relying on learned priors and general features to effectively handle unseen distortions. For instance, Zero-shot methods (e.g.[90, 91, 92, 93]) leverage pre-trained diffusion models as generative priors, seamlessly integrating degraded images as conditions into the sampling process. In the case of AiOIR, the pre-trained MPerceiver [71] demonstrates strong zero-shot and few-shot capabilities across six unseen tasks. Additionally, TAO [85] employs Test-Time Adaptation for AiOIR tasks, achieving results comparable to or better than traditional supervised methods.
Figure 3: Ten categories of the image restoration method prototype (the last nine are AiOIR methods) (a) independent encoders and decoders. (b) multiple heads and shared decoders. (c) single branch end-to-end model. (d) shared encoders and multiple decoders. (e) a reusable pretrained middle-level backbone. (f) Mixture-of-Experts architectures. (g) pretrained large vision-language models. (h) visual prompt to guide image restoration. (i) textual prompt or instructions to guide image restoration. (j) a question answering paradigm.
II-C Typical Scenarios
@@
We categorize the task of AiOIR scenario into several scenarios based on the type of images being processed: natural images, adverse weather conditions, medical images, and document images.
@@
Natural images. The degradation types commonly encountered in natural images include Gaussian noise, real noise, defocus blur, motion blur, low-light conditions, JPEG compression artifacts, mosaic effects, underwater distortions, and issues arising from under-display cameras, among others. Various AiOIR approaches have been developed to address these degradations, including techniques outlined in studies such as [63, 55, 66, 65, 61, 72, 94, 95, 73, 70, 96, 97, 98, 99]. These approaches primarily focus on resolving typical problems found in both natural and man-made scenes, enhancing the quality and usability of images affected by these common degradations.
@@
Adverse weather conditions. In more extreme cases, the images requiring restoration may be severely impacted by various adverse weather conditions, such as snowflakes or dense haze effects. These conditions lead to ill-posed inverse problems that are crucial for applications in autonomous navigation and surveillance systems. The AiOIR field has seen the emergence of multiple solutions aimed at addressing these challenges, including works like [56, 100, 64, 101, 102, 69, 67, 75, 68, 103, 104, 105]. These solutions are designed to effectively restore visibility and clarity in images affected by harsh environmental conditions.
@@
Medical images. The domain of medical imaging within AiOIR encompasses various types, including clinical Computed Tomography (CT), Magnetic Resonance Imaging (MRI), and Positron Emission Tomography (PET). Notable approaches in this area include AMIR [106] and ProCT [107]. AMIR employs a task-adaptive routing strategy that achieves state-of-the-art performance across three key medical imaging restoration tasks: MRI super-resolution, CT denoising, and PET synthesis. On the other hand, ProCT introduces innovative view-aware prompting techniques along with artifact-aware contextual learning, enabling universal incomplete-view CT reconstruction while demonstrating seamless adaptability to out-of-domain hybrid CT scenarios.
@@
Document images. Currently, there is only one notable article focused on All-in-One document image restoration, which is DocRes [108]. This work addresses a variety of tasks, including dewarping, deshadowing, appearance enhancement, deblurring, and binarization. DocRes employs a straightforward yet highly effective visual prompting method known as DTSPrompt, which effectively distinguishes between different tasks and accommodates various resolutions. This approach not only streamlines the restoration process for document images but also enhances the overall quality and readability of the restored documents.

### v2

@@ Abstract
Image restoration (IR) aims to recover high-quality images from inputs degraded by various factors such as noise, blur, compression, and adverse weather. Traditional IR methods typically focus on specific types of degradation, which limits their effectiveness in real-world scenarios with complex distortions. In response to this challenge, the all-in-one image restoration (AiOIR) paradigm has recently emerged, offering a unified framework that adeptly addresses multiple degradation types. These innovative models enhance convenience and versatility by adaptively learning degradation-specific features while simultaneously leveraging shared knowledge across diverse corruptions. In this survey, we present the first comprehensive overview of AiOIR, offering a taxonomy that organizes existing methods by architecture innovations, learning strategies, and key improvements. We systematically categorize prevailing approaches and critically assess the challenges these models encounter, proposing future research directions to propel this rapidly evolving field. Our survey begins with an introduction to the foundational concepts of AiOIR models, followed by a categorization of typical scenarios. We then highlight key architectural and algorithmic advances in AiOIR, aiming to inspire continued innovation. To facilitate rigorous evaluation of existing methods, we collate and summarize established datasets, evaluation metrics, and common experimental settings. Finally, we present an objective comparison of open-sourced methods, providing valuable insights for researchers and practitioners. This paper stands as the first comprehensive and insightful review of all-in-one image restoration. A related repository is available at https://github.com/Harbinzzy/All-in-One-Image-Restoration-Survey.
@@ I Introduction
Image processing is an integral part of low-level vision tasks, and digital image processing has evolved significantly over the past few decades, transitioning from traditional methods to advanced deep learning techniques. These methods, while effective, are limited by their inability to handle complex and varied image degradation scenarios. With the rise of deep learning, image processing has achieved remarkable results, especially driven by convolutional neural networks (CNNs) [1], Transformers [2]. In the image restoration (IR) field, single-task IR has achieved notable breakthroughs in correcting a specific type of image degradation (e.g., denoising [3, 4, 5, 6, 7, 8], dehazing [9, 10, 11, 12, 13], desnowing [14, 15, 16], deraining [17, 18, 19], deblurring [20, 21] and low-light image enhancement [22, 23, 24]). Although they have set state-of-the-art performance on restoring diverse degraded images, existing single-task IR methods generally lack the flexibility to adapt to new types of image degradation without extensive retraining (it is worth noting that some real-world IR methods [25, 26, 27, 28] are not categorized as single-task IR). To alleviate the aforementioned issues, researchers have proposed all-in-one image restoration (AiOIR), devoting significant effort to introduce various key improvements (e.g., prompt-based learning approach, mixture-of-experts structure, multimodal model and autonomous agent) beyond single-task IR.
The AiOIR approach aims to develop a model capable of addressing multiple isolated or mixed degradation tasks (e.g., denoising, deblurring, dehazing, deraining, and low-light enhancement) within a unified framework without the need for task-specific retraining. While single-task models often achieve state-of-the-art performance for specific restoration tasks, AiOIR models offer the potential for more efficient deployment and unified processing, particularly when handling unknown, mixed, or dynamically changing degradations. Despite the variety of AiOIR models, their ability to generate high-quality images is still under active exploration. Recently, pioneering researchers have been investigating ways to optimize the architectures of these models to balance the trade-offs between computational complexity and restoration quality.
@@
With the rapid development of AiOIR, numerous researchers have collected a series of datasets tailored for various all-in-one task settings, e.g., the All-Weather dataset [29] for synthetic weather-related scenarios, the WeatherStream dataset [30] for real-world weather-related scenarios, the CDD-11 dataset [31] for composite degradations, etc. Leveraging these datasets, most recent works focus on improving the representation capability of IR networks for complicated degradation through well-designed methods based on prompt learning, contrastive learning, multimodal representations, etc. Although these works have achieved superior progress in the objective quality(e.g., PSNR, NIQE [32] and FID [33]), the restored images still suffer from unsatisfactory texture generation, hindering the application of IR methods in real-world scenarios. Collectively, these methodologies represent a substantial progression in the pursuit of sophisticated, accurate, and versatile AiOIR solutions.
@@
AiOIR methods have emerged as a positive step forward, although their specific implementations and applications are still under exploration. While there are comprehensive surveys for single-task IR methods, such as image super-resolution [34, 35, 36, 37], deraining [38, 39], dehazing [40], denoising [41, 42, 43], deblurring [44, 45], low-light enhancement [46, 47, 48], and the review covering diffusion-based image restoration [49], no existing review has systematically examined the emerging field of AiOIR. To bridge this gap, we aim to provide the first comprehensive overview of AiOIR methods, shedding light on both its representative approaches and diverse improvements. Fig. 1 shows the taxonomy of this survey in a hierarchically-structured way. In the following sections, we will cover various aspects of recent advances in AiOIR:
@@
    Preliminaries (Sec. II) : We introduce the definition of AiOIR and provide a comparison with related concepts. We point out typical scenarios and task relationships.
@@
    Methodology (Sec. III) : We provide a detailed analysis of representative AiOIR networks, aiming to illustrate the prevailing methods and elucidate the different categories of methods. By analyzing state-of-the-art methods, we summarize network architectures, learning strategies, and some key improvements.
@@
    Experiments (Sec. IV) : To facilitate a reasonable and exhaustive comparison, we clarify the commonly-used datasets and experimental settings in AiOIR. Furthermore, comprehensive comparisons across different settings are provided.
@@
    Challenges and Future Directions (Sec. V) : There are still some challenges to extending AiOIR to the practical applications. To further improve the development of AiOIR, we summarize the primary challenges and propose the potential directions and trends.
@@ II Preliminaries
In this section, we first define the concept of all-in-one image restoration and compare it with related concepts. Subsequently, we provide a comprehensive review of diverse scenarios in AiOIR and the corresponding approaches. Finally, we illustrate the interactions between different IR tasks.
II-A Concept Definitions
@@
Single degradation image restoration (SDIR) focuses on recovering clean images from observations that have been corrupted by a specific type of degradation, such as noise [4, 5, 6], blur [20, 21], or rain [17, 18, 19]. These methods are typically task-specific, meaning that each model is designed for a particular degradation type. While these models demonstrate strong performance on known degradation types, they often exhibit limited generalization capability when encountering unseen degradation patterns or novel corruption intensities. Recently, some works [50, 51, 52] show certain generalizability to various degradation types with a unified framework. However, they need to train different models for different degradations, which are not all-in-one solutions as expected in practice. We refer to these methods as multiple degradation image restoration (MDIR). Subsequently, there have been recent attempts to unify multiple image restoration tasks within a single framework without the need for retraining. The seminal work [53] proposed the all-in-one weather restoration method which utilizes a multi-encoder and decoder architecture and the neural architecture search across task-specific optimized encoders. Moreover, Li et al. [54] first proposed the all-in-one blind image restoration using contrastive learning to extract degradation representations, which can recover images from multiple corruptions without any degradation information in advance. The AiOIR approach aims to recover clean images under various conditions explicitly tailored to address isolated or mixed degradation based on multi-head and multi-tail structures, priors, or pretrained models within a unified framework. It is worth noting that, some robust IR models [25, 27, 26, 28, 55, 56] are still broadly considered part of the AiOIR setting. The AiOIR models present practical advantages such as reduced storage requirements and simplified deployment. However, the principal challenge lies in developing robust architectures capable of effectively addressing diverse degradations and achieving high-quality restorations across varying conditions in real-world scenarios.
Figure 2: Milestones of AiOIR methods. We simply list their names and venues. The bar chart shows the approximate annual number of papers on AiOIR from 2020 to May 2025.
II-B Typical Scenarios
@@
We categorize AiOIR into several scenarios according to types of images processed, including natural images, images under adverse weather conditions, underwater images, medical images, document images, and hyperspectral images.
@@
Natural images. The degradation types commonly encountered in natural images include Gaussian noise, real-world noise, defocus blur, motion blur, low-light conditions, JPEG compression artifacts, mosaic effects, and issues arising from under-display cameras. Various AiOIR approaches have been developed to address these degradations, including techniques outlined in studies such as [57, 58, 59, 60, 61, 62]. These approaches primarily focus on resolving typical problems found in both natural and man-made scenes, enhancing the quality and usability of images affected by these common degradations.
@@
Adverse weather conditions. In more extreme cases, the images requiring restoration may be severely impacted by various adverse weather conditions, such as snowflakes, raindrops, and dense haze effects. These conditions lead to ill-posed inverse problems that are crucial for applications in autonomous navigation, video retrieval, and outdoor surveillance. The AiOIR field has seen the emergence of many solutions aiming at addressing these challenges, including works like [63, 29, 64, 65, 66, 67]. Notably, Liu et al. [68] proposed multi-weather nighttime image restoration. These solutions are designed to effectively restore visibility and clarity in images affected by harsh environmental conditions.
Underwater images. Underwater images suffer from wavelength-dependent light attenuation and scattering by marine microorganisms, leading to haze, color casts, low contrast, and loss of fine detail. UniUIR [56], a pioneering framework to tackle such compound distortions in an integrated manner, introduces the Mamba MoE module leveraging DepthAnythingV2 [69] to inject scene depth priors.
@@
Medical images. The domain of medical imaging within AiOIR encompasses various types, including clinical computed tomography (CT), magnetic resonance imaging (MRI), and positron emission tomography (PET). Notable approaches in this area include AMIR [70] and ProCT [71]. AMIR utilizes a task-adaptive routing strategy that achieves state-of-the-art performance across three key medical imaging restoration tasks (i.e., MRI super-resolution, CT denoising, and PET synthesis). Meanwhile, ProCT introduces an innovative view-aware prompting technique combined with artifact-aware contextual learning. This approach enables universal incomplete-view CT reconstruction and demonstrates seamless adaptability to out-of-domain hybrid CT scenarios.
@@
Document images. DocRes [72] pioneers the exploration of generalist models for all-in-one restoration of document images. This work addresses five tasks (i.e., dewarping, deshadowing, appearance enhancement, deblurring, and binarization). DocRes employs a straightforward yet highly effective visual prompting method known as DTSPrompt, which effectively distinguishes between different tasks and accommodates various resolutions. TextDoctor [73], utilizing the structure pyramid prediction and patch pyramid diffusion model, exhibits superior performance in capturing the structure of inpainted text.
Hyperspectral images. While effective for the RGB image restoration, all-in-one methods struggle with hyperspectral images (HSIs), as the complexity and variability of degradation hinder the generalization of common strategies. PromptHSI [74] integrates frequency-domain priors with the vision-language model driven prompt learning, and introduces a composite degradation dataset for all-in-one HSI restoration. MP-HSIR [75] further introduces a novel multi-prompt framework that integrates spectral, textual, and visual prompts. Unlike PromptHSI, MP-HSIR exploits complementary prompt modalities to better capture diverse degradation patterns and enhance spectral restoration.
II-C Task Interactions
While the aforementioned taxonomy outlines various scenarios for AiOIR, problems in the real world are inherently complex. It remains crucial to understand the interactions between different restoration tasks, particularly in mixed-degradation scenarios, where multiple types of corruption co-occur (e.g., low-light and blur). These task interactions can be complementary, independent, or even conflicting, significantly affecting the overall restoration quality and the design of unified models. For example:
  • Noise and blur: High ISO or low-light settings often cause both sensor noise and motion blur. Noise may obscure blur kernels, while inaccurate deblurring can amplify noise residuals [76].
  • Rain, haze, and resolution: In outdoor surveillance, rain streaks and haze not only degrade contrast but also interfere with spatial detail, complicating subsequent super-resolution.
  • Color and illumination distortions: In underwater or adverse weather conditions, color shift, low contrast, and backscatter occur simultaneously, requiring integrated correction of color, visibility, and detail [77].
These intertwined degradations present both challenges (task entanglement, ordering sensitivity, domain ambiguity) and opportunities (shared features, multi-task learning, unified representations) for AiOIR models. Understanding these interdependencies is crucial for the design of robust and generalizable AiOIR systems, especially in unconstrained or real-world environments where degradation types and combinations are unpredictable.